OpenStudy (anonymous):

To graph the equation y equals one half times x minus 3 using the slope and y-intercept, you would have to do which of the following? Graph the y-intercept, 3. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -3. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -3. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-

OpenStudy (anonymous):

2nd choice

OpenStudy (anonymous):

y=1/2x-3 y-intercept=-3 first choice is refused substitute in sec eqution when move up one y=-2 =1\2x-3 . therfore x=2 which means two moves 2 steps to the right